# Haldane model on the Sierpiński gasket

### v2

@@ IV.1 Real-Space Chern number
Here, $P$ is the projector onto the Fermi sea, and $A$, $B$, and $C$ are three touching subsets of lattice points arranged in counterclockwise order. When the bulk Chern number of the investigated system is quantised, $\nu(P)$ becomes independent of the specific choice of $A$, $B$, and $C$ in the limit where the number of sites in each subset approaches infinity. Therefore, $A$, $B$, and $C$ should be chosen such that each region is large enough to capture the system’s topology. However, they cannot span the entire system because in that case $\nu({P})=0$, per construction [59].
@@
Figure 7: The numerically calculated values of $\nu(P)$ as a function of a projector onto the states filling the lowest $\eta$ percentile of states. The red and blue values correspond to the two choices of $A$, $B$ and $C$, depicted in the insets with the respective border colour. In both cases, the same system was considered with $M=0.4t$, $\lambda=0.2$, and $\Phi=\pi/2$. The observed difference is a consequence of an undefined Chern number when the Fermi energy lies within a band.
@@
Another curious artefact is the small bump in Fig. 6(a), which can be observed at energies around $E_{f}=-t$. Further investigation of this bump reveals that this behaviour is related to the specific choice of $A$, $B$, and $C$. In Fig. 7, $\nu(P)$ is shown as a function of the projector onto all states below the Fermi energy, which corresponds to a system with filling fraction $\eta$, for two different choices of $A$, $B$, and $C$. Here, the strength of the staggered mass is also increased to $M=0.4t$ to inflate the effect.
@@
Figure 8: The interplay between $\lambda$, $M$ and $\Phi$ and the energy of the observed topological band gaps for a triangular flake [(a)-(c) and (g)-(i)] and a Sierpiński gasket [(d)-(f) and (j)-(l)]. In (a)-(f), the energy of a gap is considered, while the filling fractions are considered in (g)-(l). For each figure, one parameter is varied, while the others are set to $\lambda=0.2t$, $M=0.05t$, and $\Phi=\pi/2$. These values are marked with a black dashed line in the respective figures. Both systems have $t=1$, and have the same size. A deep pink (green) colour represents a stable value of $\tilde{\nu}(P)=1$ $(-1)$. In (a)-(c) and (g)-(i), the opening and closing of a topological gap under the conditions predicted by Haldane [4] can be observed. In (d)-(f) and (j)-(l), we observe the opening and closing of multiple topological gaps, also away from $E_{f}=0$ and $\eta=0.5$. These topological gaps have a much more complex dependence on the parameters.
@@
The two choices of $A$, $B$, and $C$ are depicted in the insets and the corresponding values of $\nu(P)$ are shown in red and blue, respectively. For most filling fractions, the red and blue graphs do not coincide, except for a small range of filling fractions around $\eta=1/2$ (marked in grey). This range corresponds to the topological states, which have an RSCN quantized exactly at $\nu(P)=1$. The RSCN is independent of the specific choice of $A$, $B$, and $C$ when the Chern number [Eq. (4)] itself is quantized. A quantized Chern number exists only as a property of a full band, but for filling fractions inside the bulk bands (outside the grey area), the Chern number is not defined and consequently $\nu(P)$ is non-physical.
@@
In a sense, Fig. 7 shows the real-space equivalent of the process of summing over the Berry curvature. How and which different choices of $A$, $B$, and $C$ compose different routes to the same topological invariant is an interesting topic of research which requires more investigation. For now, a method to determine whether a filling fraction constitutes a system with filled “bulk bands” is needed. Here, we offer two possible solutions. The first method is to consider a large set of choices for $A$, $B$, and $C$, and verify whether the values of $\nu(P)$ are in agreement, but this can be costly. A second approach is to take a numerical derivative with respect to $\eta$ and check whether the rate at which the value of $\nu(P)$ changes is small. A change of $\nu(P)$ implies that it is not yet quantized, the mode is part of a band, and the Chern number is ill-defined. However, this method might lose some “in-gap” modes close to the boundaries of topological gaps, due to the nature of a derivative. Nevertheless, since this approach requires significantly less computational time, we adopt it here. Hence, we introduce $\tilde{\nu}(P)$, which is a masked version of $\nu(P)$. This mask exponentially fades out values of the RSCN with a large numerical derivative of $\nu(P)$, with a half-life of $\nu^{\prime}_{1/2}=0.05$. Its effect is discussed in Appendix C in more detail.
@@ IV.2 Sierpiński cocoons
Similar images of a fifth-generation Sierpiński triangle are depicted in Figs. 8(d)-(f). A much richer structure is observed. When comparing the influence of complex NNN hopping strength on the topological bands in Figs. 8(a) and (d), more distinct topological gaps are observed in the case of a Sierpiński triangle. For any small value of $\lambda/t$, the model exhibits a central band gap with $\nu(P)=+1$. This is different from the triangular flakes, for which a minimal complex NNN hopping strength $\lambda=M/3\sqrt{3}$ was required to open a topological gap when $M\neq 0$. Furthermore, we also observe topological band gaps away from the Fermi level, with different Chern numbers. For larger values of $0.60<\lambda/t<$ 0.6, it becomes apparent that the central gap is formed of multiple smaller topological gaps, all with $\nu(P)=+1$. These gaps seem to close and reopen with the opposite Chern number for $10.6<\lambda/t<$ 1. This closing and reopening is a recurrent pattern, which is also recognised in the topological bands around $E\approx 1.35t$ for smaller values of $\lambda/t$, but here the topological gaps change from $-1$ to $1$. In Fig. 8(e), the Semenoff mass is shown to impact the width of these central band gaps and to lead to a slanted, diagonal band, which in not centred around at a constant value of $E$. The latter is a consequence of the imbalance in the number of sites in each sublattice and the different effect that the mass has on each sublattice. Lastly, in Fig. 8(f) we see similar tendencies for the middle band as for the topological band gaps in Fig. 8(c), and once again we observe topological gaps of different RSCN at higher energies.
In Figs. 8(g)-(l), similar figures are displayed as in Figs. 8(a)-(f), but now the filling fraction $\eta$ is considered instead of the Fermi energy. They are related by the density of states, and therefore they share many similarities. When compared to Figs. 8(a)-(f), we observe an important difference; the energy around which the topological gaps lie fluctuates, depending on the specific choice of parameters, but the filling fractions corresponding to topological gaps seem to be more stable, forming horizontal lines. In the case of a triangular flake, the topological gaps are found around half-filling, $\eta=0.5$, as expected. A similar effect is observed for the Sierpiński gasket, most prominently in Fig. 8(k), where a dependence of the energy of a gap on the parameter $M$ is converted into a constant value of $\eta$. For the Sierpiński gasket, there is a flat band around half-filling, and these states are non-topological, as best observed in Fig. 8(k). However, for any nonzero $\lambda$, the flux-induced gaps on either side of half-filling are topological, and these gaps are well-defined for all parameters.
@@ IV.3 Phase diagram
To compare the results of this work with the ones by Haldane [4] and by Li et al. [53], we present a phase diagram as a function of the staggered mass $M$ and the complex NNN hopping phase factor $\Phi$. Since the gaps are better defined in terms of a constant filling fraction $\eta$ than in terms of an energy $E$, as depicted in Figs. 8, the phase diagrams are shown for systems of constant $\eta$. For the triangular flake, we choose $\eta=0.5$ because the topological gap is predicted to be at half filling and Figs. 8(g)-(i) show a well-defined topological gap for this filling, for a wide range of both $M$ and $\Phi$. However, for the Sierpiński gasket we select $\eta=281/609\approx 0.46$. This is less than half-filling to circumvent the observed flat band, and corresponds to a flux-induced gap. Moreover, from Figs. 8(j)-(l) we know that there is a well-defined topological band gap for this fraction for a wide range of $M$ and $\Phi$. In Fig. 9, the phase diagrams are shown for (a) a triangular flake and (b) a Sierpiński gasket of the same size.
Figure 9: The phase diagram of (a) a triangular flake and (b) a fifth-generation Sierpiński gasket for a constant filling fraction $\eta$, with complex NNN hopping strength $\lambda=0.2t$, as a function of the staggered mass $M$ and the complex NNN hopping phase $\Phi$. For the triangle, $\eta=0.5$ and for the Sierpiński gasket $\eta\approx 0.46$. The colours represent the rounded RSCN. The theoretical phase boundaries given in Eq. (2) are represented in black, showing good agreement with the triangular model considered, up to some finite-size effects.
@@
Figure 11: The phase diagrams corresponding to two different filling fractions of a fifth-generation Sierpiński triangle. These two are $\eta=281/609$ and $\eta=242/609$, corresponding to a flux-induced and a fractal gap, respectively. In (a), the energy spectrum is shown for $M=0.01t$ and $\lambda=0.12t$, and the two filling fractions $\eta$ of interest are marked with blue and purple lines, together with the last filled state. In (b), the RSCNs of these gaps are shown for a set of complex NNN hopping strengths ranging from 00 to $t$. The masked RSCN $\tilde{\nu}(P)$ is represented as a colour map ranging from 1 to -1. For both filling fractions, we see the emergence of topological phases. In the top row, topological phases are observed for a wide range of parameters, showcasing intricate patterns. For the largest fractal gap, we also observe topological phases, with a similar complexity. This interesting behaviour increases with the strength of NNN hoppings and is reminiscent of Arnold’s tongues.
@@
Furthermore, the authors of Ref. [53] (and many others) quantify the topology of a fractal with the Bott index. In general, to apply the Bott index one implements (artificial) PBC, which exclude the boundaries from the calculation. This results in a new “bulk”, which can be used to predict what happens when one “cuts” this material and introduces an edge. Usually, this approximation is protected by the bulk-boundary correspondence if some part of the system is well separated from the edges. However, compared to more traditional OBC structures, the Sierpiński gasket has no well-defined bulk, and contains not only an outermost edge, but also many internal edges on the boundaries with the holes. Finding boundary conditions that eliminate these internal edges is quite complicated. However, if this is not done, these holes and their edges remain in the Bott index calculation and one should be careful to assume that this index can be used to describe the behaviour of states on these edges. We wonder whether this marker indeed describes the behaviour on the internal edges or whether it may describe the behaviour at the boundaries of a larger lattice of these fractals.
@@
In Fig. 11(a), the energy spectrum of a fifth-generation Sierpiński gasket is shown. In this spectrum, the vertical lines represent the two values of $\eta$ of interest, coloured in blue (flux-induced gap) and purple (fractal gap). All states below these lines are filled and the last filled state is marked in the corresponding colour. In Fig. 11(b), phase diagrams in terms of $\Phi$ and $M/t$ are shown for both of these gaps for six different values of the complex NNN hopping amplitude $\lambda$. Here, the background colours (blue and purple) indicate to which filling fraction $\eta$ the phase diagram belongs. The pink and green colours indicate the masked RSCN $\tilde{\nu}(P)$ of a fifth-generation Sierpiński gasket with staggered mass $M$ and complex NNN hopping phase $\Phi$ for the corresponding filling fraction. As expected, when the NNN hopping strength $\lambda=0$, there are no topological phases, for both values of $\eta$. However, for $\lambda=0.2$ and $\eta\approx 0.46$ (blue), a wide range of topological states is observed, which are anti-symmetric around $\Phi=0$. As $\lambda$ is increased further, (inverted) conical shapes form around $\Phi=0$ and $\Phi=\pm\pi$. Furthermore, when $\lambda\geq 0.6$, two droplets start to emerge around $\Phi=\pm\pi/2$ and $M=0$, with opposite RSCN compared to their surroundings.
For the largest fractal gap, which corresponds to $\eta\approx 0.4$ (purple), topological regions induced by $M$ and $\Phi$ are still observed, but now they are more restricted. Increasing the value of $\lambda$ results in a rising level of complex behaviour, reminiscent of Arnold’s tongues [62]. Such Arnold tongues are also observed in the fractal seen in the parameter space of dynamical systems [63]. Here, interesting and complex behaviour seems to grow from the centre outwards on top of some conic shape. This conic shape reaches lower values of $M$ by increasing $\lambda$. At values of $\lambda$ between $0.8t$ and $t$, a droplet emerges at $\Phi=\pm\pi$ and $M=0$. At these filling fractions, the sinusoidal shape of the original Haldane model is completely lost, and the fractality of this system seems to induce much more intricate behaviour. In Appendix D, the influence of the chosen generation of the fractal is investigated. There, we present the phase diagram of a third, fourth, fifth, and sixth generation of the Sierpiński gaskets with a strong complex NNN hopping, $\lambda=0.8t$.
@@
Figure 12: The effect of disorder on six edge states (coloured) shown in Fig. 5 and three additional bulk states (grey). Here, some arbitrary on-site disorder between $-\Delta$ and $\Delta$ is introduced. The disorder parameter $\Delta$ is $0.02\lambda$, and $0.2\lambda$ for (a) and (b), respectively.

### v3

@@ IV.1 Real-Space Chern number
Here, $P$ is the projector onto the Fermi sea, and $A$, $B$, and $C$ are three touching subsets of lattice points arranged in counterclockwise order. Therefore, $P_{jk}$ is the ${j,k}$th element of the projector $P$ onto the occupied states at a given Fermi energy, when $P$ is written in the basis of local atomic orbitals. When the bulk Chern number of the investigated system is quantized, $\nu(P)$ becomes independent of the specific choice of $A$, $B$, and $C$ in the limit where the number of sites in each subset approaches infinity. Therefore, $A$, $B$, and $C$ should be chosen such that each region is large enough to capture the system’s topology. However, they cannot span the entire system because in that case $\nu({P})=0$, per construction [71].
@@
Figure 7: The numerically calculated values of $\nu(P)$ as a function of a projector onto the states filling the lowest $\eta$ percentile of states. The red and blue values correspond to the two choices of $A$, $B$ and $C$, depicted in the insets with the respective border color. In both cases, the same system was considered with $M=0.4t$, $\lambda=0.2$, and $\Phi=\pi/2$. The observed difference is a consequence of an undefined Chern number when the Fermi energy lies within a band.
@@
Another curious artefact is the small bump in Fig. 6(a), which can be observed at energies around $E_{f}=-t$. Further investigation of this bump reveals that this behavior is related to the specific choice of $A$, $B$, and $C$. In Fig. 7, $\nu(P)$ is shown as a function of the projector onto all states below the Fermi energy, which corresponds to a system with filling fraction $\eta$, for two different choices of $A$, $B$, and $C$. Here, the strength of the staggered mass is also increased to $M=0.4t$ to inflate the effect.
@@
Figure 8: The interplay between $\lambda$, $M$ and $\Phi$ and the energy of the observed topological band gaps for a triangular flake [(a)-(c) and (g)-(i)] and a Sierpiński gasket [(d)-(f) and (j)-(l)]. In (a)-(f), the energy of a gap is considered, while the filling fractions are considered in (g)-(l). For each figure, one parameter is varied, while the others are set to $\lambda=0.2t$, $M=0.05t$, and $\Phi=\pi/2$. These values are marked with a black dashed line in the respective figures. Both systems have $t=1$, and have the same size. A deep pink (green) color represents a stable value of $\tilde{\nu}(P)=1$ $(-1)$. In (a)-(c) and (g)-(i), the opening and closing of a topological gap under the conditions predicted by Haldane [4] can be observed. In (d)-(f) and (j)-(l), we observe the opening and closing of multiple topological gaps, also away from $E_{f}=0$ and $\eta=0.5$. These topological gaps have a much more complex dependence on the parameters.
@@
The two choices of $A$, $B$, and $C$ are depicted in the insets and the corresponding values of $\nu(P)$ are shown in red and blue, respectively. For most filling fractions, the red and blue graphs do not coincide, except for a small range of filling fractions around $\eta=1/2$ (marked in gray). This range corresponds to the topological states, which have an RSCN quantized exactly at $\nu(P)=1$. The RSCN is independent of the specific choice of $A$, $B$, and $C$ when the Chern number [Eq. (4)] itself is quantized. A quantized Chern number exists only as a property of a full band, but for filling fractions inside the bulk bands (outside the gray area), the Chern number is not defined and consequently $\nu(P)$ is non-physical.
@@
In a sense, Fig. 7 shows the real-space equivalent of the process of summing over the Berry curvature. How and which different choices of $A$, $B$, and $C$ compose different routes to the same topological invariant is an interesting topic of research which requires more investigation. For now, a method to determine whether a filling fraction constitutes a system with filled “bulk bands” is needed. Here, we offer two possible solutions. The first method is to consider a large set of choices for $A$, $B$, and $C$, and verify whether the values of $\nu(P)$ are in agreement, but this can be costly. A second approach is to take a numerical derivative with respect to $\eta$ and check whether the rate at which the value of $\nu(P)$ changes is small. A change of $\nu(P)$ implies that it is not yet quantized, the mode is part of a band, and the Chern number is ill-defined. However, this method might lose some “in-gap” modes close to the boundaries of topological gaps, due to the nature of a derivative. Nevertheless, since this approach requires significantly less computational time, we adopt it here. Hence, we introduce $\tilde{\nu}(P)$, which is a masked version of $\nu(P)$. This mask exponentially fades out values of the RSCN with a large numerical derivative of $\nu(P)$, with a half-life of $\nu^{\prime}_{1/2}=0.05$. Its effect is discussed in Appendix D in more detail.
@@ IV.2 Sierpiński cocoons
The results of a similar procedure on a fifth-generation Sierpiński triangle are depicted in Figs. 8(d)-(f). A much richer structure is observed. When comparing the influence of complex NNN hopping strength on the topological bands in Figs. 8(a) and (d), more distinct topological gaps are observed in the case of a Sierpiński triangle. For any small value of $\lambda/t$, the model exhibits a central band gap with $\tilde{\nu}(P)=+1$. This is different from the triangular flakes, for which a minimal complex NNN hopping strength $\lambda=M/3\sqrt{3}$ was required to open a topological gap when $M\neq 0$. For larger values of $0.60<\lambda/t<$ 0.6, it becomes apparent that the central gap is formed of multiple smaller topological gaps, all with $\tilde{\nu}(P)=+1$. These gaps seem to close and reopen with the opposite Chern number for $10.6<\lambda/t<$ 1. This is in stark contrast with the results on a triangular flake, where just a single topological gap opens around $E_{f}=0$, and no phase transition is observed. Furthermore, there are also topological band gaps away from the Fermi level, with different Chern numbers. The closing and reopening of topological gaps is a recurrent pattern, also visible in the topological bands around $E\approx 1.35t$ for smaller values of $\lambda/t$. However, in this case the topological gaps change from $\tilde{\nu}(P)=-1$ to $\tilde{\nu}(P)=1$. In Fig. 8(e), the Semenoff mass is shown to impact the width of these central band gaps and to lead to a slanted, diagonal band. The latter is a consequence of the imbalance in the number of sites in each sublattice and the different effect that the mass has on each sublattice. Again, topological gaps appear at energies away from $E_{f}=0$. The structure acquires the shape of a letter ‘$Z$’, pierced by multiple gaps. Lastly, in Fig. 8(f) the more dense cocoon shape from Fig. 8(c) becomes smaller and consists of multiple distinct gaps. Once again, we observe more bands and topological gaps of different RSCN at higher energies.
In Figs. 8(g)-(l), similar figures are displayed as in Figs. 8(a)-(f), but now the filling fraction $\eta$ is considered instead of the Fermi energy. They are related by the density of states, and therefore share many similarities. When compared to Figs. 8(a)-(f), we observe an important difference; the energy around which the topological gaps lie fluctuates, depending on the specific choice of parameters, but the filling fractions corresponding to topological gaps seem to be more stable, forming horizontal lines. In the case of a triangular flake, the topological gaps are found around half-filling, $\eta=0.5$, as expected. A similar effect is observed for the Sierpiński gasket, most prominently in Fig. 8(k), where a dependence of the energy of a gap on the parameter $M$ is converted into a constant value of $\eta$. For the Sierpiński gasket, there is a flat band around half-filling, and these states are non-topological, as best observed in Fig. 8(k). However, for any nonzero $\lambda$, the flux-induced gaps on either side of half-filling are topological, and these gaps are well-defined for all parameters.
@@ IV.3 Phase diagram
To compare the results of this work with the ones by Haldane [4] and by Li et al. [64], we present a phase diagram as a function of the staggered mass $M$ and the complex NNN hopping phase factor $\Phi$. Since the gaps are better defined in terms of a constant filling fraction $\eta$ than in terms of an energy $E$, as depicted in Figs. 8, the phase diagrams are shown for systems of constant $\eta$. For the triangular flake, we choose $\eta=0.5$ because the topological gap is predicted to be at half filling and Figs. 8(g)-(i) show a well-defined topological gap for this filling, for a wide range of both $M$ and $\Phi$. However, for the Sierpiński gasket we select $\eta=281/609\approx 0.46$. This is less than half-filling to circumvent the observed flat band. Moreover, from Figs. 8(j)-(l) we know that there is a well-defined topological band gap for this fraction for a wide range of $M$ and $\Phi$. We have chosen these values of $\eta$ because both of them correspond to a gap opened by a finite complex NNN hopping. A more direct comparison with the same values of $\eta$ would be desirable but the fractal nature of the Sierpiński gasket makes such a comparison impossible because of the emergence of a central flat band consisting of compact localized states [57]. In Fig. 9, the phase diagrams are shown for (a) a triangular flake and (b) a Sierpiński gasket of the same size.
Figure 9: The phase diagram of (a) a triangular flake and (b) a fifth-generation Sierpiński gasket for a constant filling fraction $\eta$, with complex NNN hopping strength $\lambda=0.2t$, as a function of the staggered mass $M$ and the complex NNN hopping phase $\Phi$. For the triangle, $\eta=0.5$ and for the Sierpiński gasket $\eta\approx 0.46$. The colors represent the rounded RSCN. The theoretical phase boundaries given in Eq. (2) are represented in black, showing good agreement with the triangular model considered, up to some finite-size effects.
@@
Figure 11: The phase diagrams corresponding to two different filling fractions of a fifth-generation Sierpiński triangle. These two are $\eta=281/609$ and $\eta=242/609$, corresponding to a flux-induced and a fractal gap, respectively. In (a), the energy spectrum is shown for $M=0.01t$ and $\lambda=0.12t$, and the two filling fractions $\eta$ of interest are marked with blue and purple lines, together with the last filled state. In (b), the RSCNs of these gaps are shown for a set of complex NNN hopping strengths ranging from 00 to $t$. The masked RSCN $\tilde{\nu}(P)$ is represented as a color map ranging from 1 to -1. For both filling fractions, we see the emergence of topological phases. In the top row, topological phases are observed for a wide range of parameters, showcasing intricate patterns. For the largest fractal gap, we also observe topological phases, with a similar complexity. This interesting behavior increases with the strength of NNN hoppings and is reminiscent of Arnold’s tongues.
@@
Furthermore, the authors of Ref. [64] (and many others) quantify the topology of a fractal with the Bott index. The Bott index is a global quantity describing the bulk. In general, to apply the Bott index one implements (artificial) PBC, and one then performs a global calculation, after which the behavior at the edge is inferred via the bulk-boundary correspondence. However, compared to more traditional OBC structures, the Sierpiński gasket has no well-defined bulk because it contains not only an outermost edge, but also many internal edges on the boundaries of the voids. Finding boundary conditions that eliminate these internal edges is quite complicated. If this is not done, these holes and their edges remain in the Bott index calculation. Furthermore, applying periodic boundary conditions to a Sierpinksi gasket introduces a new length scale, which breaks the self-similarity and the dimensionality on scales larger than this length scale. This change of the dimensionality could have non-trivial effects. Taken together, it is questionable whether this marker indeed describes the behavior on the internal edges or at the boundaries of a larger lattice of these fractals. On the other hand, the RSCN, provides global information based on a local calculation, which does not require the notion of a well-separated bulk. Consequently, the RSCN method only requires the local wavefunctions in real space for the calculation of the topological invariant, and is more adequate than the Bott index in the case of fractals.
@@
In Fig. 11(a), the energy spectrum of a fifth-generation Sierpiński gasket is shown. In this spectrum, the vertical lines represent the two values of $\eta$ of interest, colored in blue (flux-induced gap) and purple (fractal gap). All states below these lines are filled and the last filled state is marked in the corresponding color. In Fig. 11(b), phase diagrams in terms of $\Phi$ and $M/t$ are shown for both of these gaps for six different values of the complex NNN hopping amplitude $\lambda$. Here, the background colors (blue and purple) indicate to which filling fraction $\eta$ the phase diagram belongs. The pink and green colors indicate the masked RSCN $\tilde{\nu}(P)$ of a fifth-generation Sierpiński gasket with staggered mass $M$ and complex NNN hopping phase $\Phi$ for the corresponding filling fraction. As expected, when the NNN hopping strength $\lambda=0$, there are no topological phases, for both values of $\eta$. However, for $\lambda=0.2$ and $\eta\approx 0.46$ (blue), a wide range of topological states is observed, which are anti-symmetric around $\Phi=0$. As $\lambda$ is increased further, (inverted) conical shapes form around $\Phi=0$ and $\Phi=\pm\pi$. Furthermore, when $\lambda\geq 0.6$, two droplets start to emerge around $\Phi=\pm\pi/2$ and $M=0$, with opposite RSCN compared to their surroundings.
For the largest fractal gap, which corresponds to $\eta\approx 0.4$ (purple), topological regions induced by $M$ and $\Phi$ are still observed, but now they are more restricted. Increasing the value of $\lambda$ results in a rising level of complex behavior, reminiscent of Arnold’s tongues [74]. Such Arnold tongues are also observed in the fractal seen in the parameter space of dynamical systems [75]. Here, interesting and complex behavior seems to grow from the center outwards on top of some conic shape. This conic shape reaches lower values of $M$ by increasing $\lambda$. At values of $\lambda$ between $0.8t$ and $t$, a droplet emerges at $\Phi=\pm\pi$ and $M=0$. At these filling fractions, the sinusoidal shape of the original Haldane model is completely lost, and the fractality of this system seems to induce much more intricate behavior. In Appendix E, the influence of the chosen generation of the fractal is investigated. There, we present the phase diagram of a third, fourth, fifth, and sixth generation of the Sierpiński gaskets with a strong complex NNN hopping, $\lambda=0.8t$.
@@
Figure 12: The effect of disorder on six edge states (colored) shown in Fig. 5 and three additional bulk states (gray). Here, some arbitrary on-site disorder between $-\Delta$ and $\Delta$ is introduced. The disorder parameter $\Delta$ is $0.02\lambda$, and $0.2\lambda$ for (a) and (b), respectively.